OpenStudy (anonymous):

Which statements are true? Choose all answers that are correct. A. –13 + 25 = 10 + (–22) B. 20 + (–6) > 31 + (–19) C. 8 + (–12) < –17 + 14 D. –11 + (–15) > 6 + (–29) i already have them figured out but i want to know if im right my answers are 14>12. -26>-23 b and d so am i correct

OpenStudy (anonymous):

@Loser66 @Gabylovesyou

OpenStudy (anonymous):

@iGreen @Starr_DynastyT

OpenStudy (igreen):

–13 + 25 = 10 + (–22) -13 + 25 = 10 - 22 12 = -12 INCORRECT

OpenStudy (igreen):

20 + (–6) > 31 + (–19) 20 - 6 > 31 - 19 14 > 31 - 19 14 > 12 CORRECT

OpenStudy (igreen):

8 + (–12) < –17 + 14 8 - 12 < -17 + 14 -4 < -17 + 14 -4 < -3 CORRECT

OpenStudy (igreen):

–11 + (–15) > 6 + (–29) -11 - 15 > 6 - 29 -26 > 6 - 29 -26 > -23 INCORRECT

OpenStudy (igreen):

So your answers are B and C.

OpenStudy (anonymous):

kk